4. Here is Benders total value schedule for cigars/day. Q TV 1 10 2 18 3 24 4 28 5 30 6 31 7 31 a. Derive a demand curve for cigars for Bender. (Il get you started) P 246 810 12 b. If the price of cigars is S8, how many does Bender buy, what is his total expenditure and consumer surplus? c. If Bender had the opportunity to join a cigar-of-the-month club which allowed him to buy cigars for $4 each, how much would he be willing to pay to join this club?
Answer #1

Calculate Marginal Value of cigars as below: MV-TVn-TVn-1 10 10 18 24 28 30 31 31 4 4 The demand curve can be derived from the MV schedule as below 4 4 10 12 Price 12 5 Output

Total Q at P- $8 will be 2 This is because Marginal value is $8 when Q 2 and while formulating demand schedule, we look at the points where P-MV Total expenditure-Po-8(2) = $16 Consumer surplus (1/2)(12-8)2) 4 At p S4, total Q 4 units Benefits accrued = (6-4) + (8-4) + (10-4) + (12-4) = $20 Thus, Bender will pay anything less than $20 to join the club